# Anytime-valid t-tests and confidence sequences for Gaussian means with unknown variance

### v4

@@ 6.2 The Surprising $\alpha$-Suboptimality of the Classical t-Test
E-values for $\mathcal{N}_{\mu}$ via one single observation $X_{1}$ can be constructed by methods other than (57). For example, in anticipation of the reverse information projection to be introduced in the next subsection, and fully formulated in Section C.4 by Proposition C.3, the following likelihood ratio,
$B:=\frac{\mathrm{d}N_{0,1}}{\mathrm{d}N_{\mu,\mu^{2}+1}}(X_{1})=\sqrt{\mu^{2}+1}\exp\left(\frac{(X_{1}-\mu)^{2}}{2(\mu^{2}+1)}-\frac{X_{1}^{2}}{2}\right),$ (62)
is an e-value for the entire set $\mathcal{N}_{\mu}$, not only for $N_{\mu,\mu^{2}+1}$, meaning that $\mathrm{E}_{P}[B]\leqslant 1$ for any $P\in\mathcal{N}_{\mu}$. (62), however, does not lead to a closed-form CI via inversion.
@@
(60) is an instance of “one-observation confidence intervals” for $\mu$ over $\mathcal{N}$, a classical topic studied by numerous authors including Charles M. Stein, Abbott and Rosenblatt [1962], and Portnoy [2019] who studies the lower bound in this problem [^2] The paper by Portnoy [2019] contains an important typo: the two $\Phi(\lambda(1+\tfrac{1}{c_{1}}))$ terms should be $\Phi(\lambda(1-\tfrac{1}{c_{1}}))$ instead, in the main result Theorem 1.1.. These constructions typically look like $[X_{1}\pm c_{\alpha}|X_{1}|]$, where, for example, $c_{0.05}\approx 10$, and are often tighter than (60). Our CI (60), however, enjoys a simple closed-form expression that requires no numerical inversion, and is proved directly by the e-value (57) to provide intuition. All these CIs including our (60), we note, contain 0 (or another fixed constant by shifting) regardless of the value of $X_{1}$. These prior results, together with our (60), showcase the lesser-known suboptimality of the classical t-test CI under small samples; in particular, its unnecessary vacuity when $n=1$. On a separate note, our one-observation CI (60) can be randomly and losslessly tightened due to a recent randomized Markov’s inequality by Ramdas and Manole [2023, Theorem 1.2]. This randomization halves the expected width of the CI. To elaborate, the randomized CI has the $1/\alpha$ term in (60) replaced by a $U/\alpha$, where $U$ is an independent uniform random variable on $[0,1]$, which halves the expected width without changing its validity.
@@ 7 Concluding Remarks
While we only focus on the class of one-dimensional normal distributions $\mathcal{N}$ in this paper, we expect similar techniques may lead to sequential Hotelling’s T-tests (with test processes and confidence sphere sequences [Chugg et al., 2023]) for the class of multivariate normal distributions. A slightly different way to generalize in dimensionality is the sequential F-tests for linear models, on which the most recent and comprehensive work by Lindon et al. [2022] only covers one of the numerous methods in our paper (namely, Gaussian mixture of scale invariant likelihood ratios). For potential nonparametric extensions of our work beyond normal distributions, one may need to appeal to asymptotic sequential tests and confidence sequences, proposed recently by Waudby-Smith et al. [2021] — whose Z-test-like results, however, arguably stand as nonparametric, asymptotic extensions of our t-tests already. We leave these topics for future authors to study in depth.
@@ Appendix A Additional Sequential t-Tests in the Literature
In this section, we review the method by Jourdan et al. [2023, Appendix E], who first derive a confidence sequence for the variance $\sigma^{2}$ based on the sample variance $v_{n}$ over $\mathcal{N}$ (using an example by Howard et al. [2021]), and then use a standard Z-confidence sequence via the method of likelihood mixture (e.g. Equation 10 of Howard et al. [2021]) that requires knowing the true variance $\sigma^{2}$. They then use a union bound to combine the two to obtain a rectangular confidence sequence for the pair $(\mu,\sigma^{2})$ over $\mathcal{N}$. A tighter confidence sequence can be obtained, we remark, if only the estimation of $\mu$ is of interest.
@@
We also let $\zeta$ be the Riemann Zeta function, which naturally arises in stitching [Howard et al., 2021], a technique for obtaining law-of-the-iterated-logarithm rate confidence sequences. Recall that we denote by $v_{n}$ the sample variance without Bessel correction,
$v_{n}=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\widehat{\mu}_{n})^{2}.$ (68)
@@
$\left[0,\frac{v_{n}}{\overline{W_{0}}\left(1+\frac{2(1+\eta)}{n-1}\left(\log\frac{\zeta(s)}{\alpha}+s\log\left(1+\frac{\log(n-1)}{\log(1+\eta)}\right)\right)\right)-\frac{1}{n-1}}\right]$ (69)
@@
$\left[\widehat{\mu}_{n}\pm\sqrt{\frac{v_{n}}{n}\cdot\frac{\overline{W}_{-1}\left(1+2\log(2/\alpha)+2\log\zeta(s)+2s(1-\log(2s))+2s\log(2s+\log n)\right)}{\overline{W_{0}}\left(1+\frac{2(1+\eta)}{n-1}\left(\log\frac{2\zeta(s)}{\alpha}+s\log\left(1+\frac{\log(n-1)}{\log(1+\eta)}\right)\right)\right)-\frac{1}{n-1}}}\right]$ (71)
@@ A.2 A Sequential Test for Median by Howard and Ramdas [2022]
Again, as is shown in Figures 1 and 2, this asymptotically optimal rate does not translate into practical advantages for sample size $<10000$.
@@ C.2 Proofs for Universal Inference t-Tests
Prood of Proposition 3.3.
Note that
@@ C.3 Proofs for Scale Invariant t-Tests
Now recall $V_{n}:(x_{1},x_{2},\dots,x_{n})\mapsto(x_{1}/|x_{1}|,x_{2}/|x_{1}|,\dots,x_{n}/|x_{1}|)$. Let $Q_{n}^{*+}$ be the push-forward measure of $Q_{n}^{+}$ by $V_{n}$, a measure on $\{1\}\times\mathbb{R}^{n-1}$ (which is the conditional distribution of $(X_{1}/|X_{1}|,\dots,X_{n}/|X_{1}|)$ given $\{X_{1}>0\}$). Suppose the set $V_{n}(A)=\{1\}\times B$. We have the following change of variables (letting $(x_{1}^{*},\dots,x_{n}^{*})=V_{n}(x_{1},\dots,x_{n})$, so $x_{1}^{*}=1$ below),
@@
First, under $N_{\mu,\sigma^{2}}$, note that $T_{n-1}=\sqrt{n-1}\frac{S_{n}-n\mu}{\sqrt{nV_{n}-S_{n}^{2}}}$ and $v_{n}=V_{n}/n-S_{n}^{2}/n^{2}$. We replace $\frac{T_{n-1}}{\sqrt{n-1}}$ by $\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu_{1}}{\sqrt{v_{n}}}$ in (34) to have
$H_{n}=\sqrt{\frac{2\pi}{n}}\left(1+\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{\sqrt{v_{n}}}\right)^{2}\right)^{n/2},$ (179)
$\frac{\log H_{n}}{n}=\frac{\log(2\pi/n)}{2n}+\frac{1}{2}\log\left(1+\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{\sqrt{v_{n}}}\right)^{2}\right).$ (180)
It is well known $v_{n}$ converges to $\sigma^{2}$ almost surely. Therefore $\frac{T_{n-1}}{\sqrt{n-1}}=\frac{\widehat{\mu}_{n}-\mu_{1}}{\sqrt{v_{n}}}$ converges almost surely to 0. These imply that $\frac{\log H_{n}}{n}$ converges almost surely to $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$.
@@
First, under $N_{\mu,\sigma^{2}}$, we replace in (42) $\frac{T_{n-1}}{\sqrt{n-1}}$ by $\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{\sqrt{v_{n}}}$,
$G_{n}^{(c)}=\sqrt{\frac{c^{2}}{n+c^{2}}}\left(1+\frac{n}{\frac{n+c^{2}}{\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{\sqrt{v_{n}}}\right)^{2}}+{c^{2}}}\right)^{n/2},$ (182)
$\frac{\log G_{n}^{(c)}}{n}=\frac{\log(\frac{c^{2}}{n+c^{2}})}{2n}+\frac{1}{2}\log\left(1+\frac{n}{\frac{n+c^{2}}{\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{\sqrt{v_{n}}}\right)^{2}}+{c^{2}}}\right).$ (183)
As in the proof of Proposition 4.8, $\frac{T_{n-1}}{\sqrt{n-1}}$ converges to 0 and $v_{n}$ to $\sigma^{2}$ almost surely. Therefore $\frac{\log G_{n}^{(c)}}{n}$ converges a.s. to $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$.

### v5

@@ 6.2 The Surprising $\alpha$-Suboptimality of the Classical t-Test
E-values for $\mathcal{N}_{\mu=\mu_{0}}$ via one single observation $X_{1}$ can be constructed by methods other than (59). For example, in anticipation of the reverse information projection to be introduced in the next subsection, and fully formulated in Section C.4 by Proposition C.3, the following likelihood ratio,
$B:=\frac{\mathrm{d}N_{0,1}}{\mathrm{d}N_{\mu_{0},\mu^{2}_{0}+1}}(X_{1})=\sqrt{\mu^{2}_{0}+1}\exp\left(\frac{(X_{1}-\mu_{0})^{2}}{2(\mu_{0}^{2}+1)}-\frac{X_{1}^{2}}{2}\right),$ (64)
is an e-value for the entire set $\mathcal{N}_{\mu=\mu_{0}}$, not only for $N_{\mu_{0},\mu^{2}_{0}+1}$, meaning that $\mathrm{E}_{P}[B]\leqslant 1$ for any $P\in\mathcal{N}_{\mu_{0}}$. (64), however, does not lead to a closed-form CI via inversion.
@@
(62) is an instance of “one-observation confidence intervals” for $\mu$ over $\mathcal{N}$, a classical topic studied by numerous authors including Charles M. Stein, Abbott and Rosenblatt [1962], and Portnoy [2019] who studies the lower bound in this problem [^4] The paper by Portnoy [2019] contains an important typo: the two $\Phi(\lambda(1+\tfrac{1}{c_{1}}))$ terms should be $\Phi(\lambda(1-\tfrac{1}{c_{1}}))$ instead, in the main result Theorem 1.1.. These constructions typically look like $[X_{1}\pm c_{\alpha}|X_{1}|]$, where, for example, $c_{0.05}\approx 10$, and are often tighter than (62). Our CI (62), however, enjoys a simple closed-form expression that requires no numerical inversion, and is proved directly by the e-value (59) to provide intuition. All these CIs including our (62), we note, contain 0 (or another fixed constant by shifting) regardless of the value of $X_{1}$. These prior results, together with our (62), showcase the lesser-known suboptimality of the classical t-test CI under small samples; in particular, its unnecessary vacuity when $n=1$. On a separate note, our one-observation CI (62) can be randomly and losslessly tightened due to a recent randomized Markov’s inequality by Ramdas and Manole [2024, Theorem 1.2]. This randomization halves the expected width of the CI. To elaborate, the randomized CI has the $1/\alpha$ term in (62) replaced by a $U/\alpha$, where $U$ is an independent uniform random variable on $[0,1]$, which halves the expected width without changing its validity.
@@ 7 Concluding Remarks
While we only focus on the class of one-dimensional normal distributions $\mathcal{N}$ in this paper, we expect similar techniques may lead to sequential Hotelling’s T-tests (with test processes and confidence sphere sequences [Chugg et al., 2023]) for the class of multivariate normal distributions. A slightly different way to generalize in dimensionality is the sequential F-tests for linear models, on which the most recent and comprehensive work by Lindon et al. [2022] only covers one of the numerous methods in our paper (namely, Gaussian mixture of scale invariant likelihood ratios). We leave these topics for future authors to study in depth.
@@ Appendix A Additional Sequential t-Tests in the Literature
In this section, we review the method by Jourdan et al. [2023, Appendix E], who first derive a confidence sequence for the variance $\sigma^{2}$ based on the sample variance $s_{n}^{2}$ over $\mathcal{N}$ (using an example by Howard et al. [2021]), and then use a standard Z-confidence sequence via the method of likelihood mixture (e.g. Equation 10 of Howard et al. [2021]) that requires knowing the true variance $\sigma^{2}$. They then use a union bound to combine the two to obtain a rectangular confidence sequence for the pair $(\mu,\sigma^{2})$ over $\mathcal{N}$. A tighter confidence sequence can be obtained, we remark, if only the estimation of $\mu$ is of interest.
@@
We also let $\zeta$ be the Riemann Zeta function, which naturally arises in stitching [Howard et al., 2021], a technique for obtaining law-of-the-iterated-logarithm rate confidence sequences. Recall that we denote by $s_{n}^{2}$ the sample variance without Bessel correction,
$s_{n}^{2}=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\widehat{\mu}_{n})^{2}.$ (70)
@@
$\left[0,\frac{s_{n}^{2}}{\overline{W_{0}}\left(1+\frac{2(1+\eta)}{n-1}\left(\log\frac{\zeta(s)}{\alpha}+s\log\left(1+\frac{\log(n-1)}{\log(1+\eta)}\right)\right)\right)-\frac{1}{n-1}}\right]$ (71)
@@
$\left[\widehat{\mu}_{n}\pm\sqrt{\frac{s_{n}^{2}}{n}\cdot\frac{\overline{W}_{-1}\left(1+2\log(2/\alpha)+2\log\zeta(s)+2s(1-\log(2s))+2s\log(2s+\log n)\right)}{\overline{W_{0}}\left(1+\frac{2(1+\eta)}{n-1}\left(\log\frac{2\zeta(s)}{\alpha}+s\log\left(1+\frac{\log(n-1)}{\log(1+\eta)}\right)\right)\right)-\frac{1}{n-1}}}\right]$ (73)
@@ A.2 A Sequential Test for Median by Howard and Ramdas [2022]
Again, as is shown in Figures 3 and 4, this asymptotically optimal rate does not translate into practical advantages for sample size $<10000$.
@@ C.2 Proofs for Universal Inference t-Tests
Proof of Proposition 3.3.
Let $P$ be a distribution with mean $\mu$ and variance $\sigma^{2}$. Note that
@@ C.3 Proofs for Scale Invariant t-Tests
Now recall $\varphi_{n}:(x_{1},x_{2},\dots,x_{n})\mapsto(x_{1}/|x_{1}|,x_{2}/|x_{1}|,\dots,x_{n}/|x_{1}|)$. Let $Q_{n}^{*+}$ be the push-forward measure of $Q_{n}^{+}$ by $\varphi_{n}$, a measure on $\{1\}\times\mathbb{R}^{n-1}$ (which is the conditional distribution of $(X_{1}/|X_{1}|,\dots,X_{n}/|X_{1}|)$ given $\{X_{1}>0\}$). Suppose the set $V_{n}(A)=\{1\}\times B$. We have the following change of variables (letting $(x_{1}^{*},\dots,x_{n}^{*})=V_{n}(x_{1},\dots,x_{n})$, so $x_{1}^{*}=1$ below),
@@
First, under any distribution with mean $\mu$ and variance $\sigma^{2}$, note that $T_{n-1}=\sqrt{n-1}\frac{S_{n}-n\mu}{\sqrt{nV_{n}-S_{n}^{2}}}$ and $s_{n}^{2}=V_{n}/n-S_{n}^{2}/n^{2}$. We replace $\frac{T_{n-1}}{\sqrt{n-1}}$ by $\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{{s_{n}}}$ in (33) to have
$H_{n}=\sqrt{\frac{2\pi}{n}}\left(1+\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{{s_{n}}}\right)^{2}\right)^{n/2},$ (181)
$\frac{\log H_{n}}{n}=\frac{\log(2\pi/n)}{2n}+\frac{1}{2}\log\left(1+\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{{s_{n}}}\right)^{2}\right).$ (182)
It is well known $s^{2}_{n}$ converges to $\sigma^{2}$ almost surely. Therefore $\frac{T_{n-1}}{\sqrt{n-1}}=\frac{\widehat{\mu}_{n}-\mu}{{s_{n}}}$ converges almost surely to 0. These imply that $\frac{\log H_{n}}{n}$ converges almost surely to $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$.
@@
First, under any distribution with mean $\mu$ and variance $\sigma^{2}$, we replace in (41) $\frac{T_{n-1}}{\sqrt{n-1}}$ by $\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{s_{n}}$,
$G_{n}^{(c)}=\sqrt{\frac{c^{2}}{n+c^{2}}}\left(1+\frac{n}{\frac{n+c^{2}}{\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{s_{n}}\right)^{2}}+{c^{2}}}\right)^{n/2},$ (184)
$\frac{\log G_{n}^{(c)}}{n}=\frac{\log(\frac{c^{2}}{n+c^{2}})}{2n}+\frac{1}{2}\log\left(1+\frac{n}{\frac{n+c^{2}}{\left(\frac{T_{n-1}}{\sqrt{n-1}}+\frac{\mu}{s_{n}}\right)^{2}}+{c^{2}}}\right).$ (185)
As in the proof of Proposition 4.8, $\frac{T_{n-1}}{\sqrt{n-1}}$ converges to 0 and $s_{n}^{2}$ to $\sigma^{2}$ almost surely. Therefore $\frac{\log G_{n}^{(c)}}{n}$ converges a.s. to $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$.
@@
Proof of Theorem 4.12.
Recall that Theorem 4.9 states that, letting
$\operatorname{radius}_{n}=\sqrt{\underbrace{\frac{(n+c^{2})\left(1-\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}\right)}{\left\{\left(\frac{\alpha^{2}c^{2}}{n+c^{2}}\right)^{1/n}(n+c^{2})-c^{2}\right\}\vee 0}}_{\rho^{2}(n,c,\alpha)}\ \cdot\ \underbrace{\left(\overline{X^{2}_{n}}-\widehat{\mu}_{n}^{2}\right)}_{s_{n}^{2}(X)}},$ (199)
then the intervals
$\left[\widehat{\mu}_{n}(X)\pm\operatorname{radius}_{n}\right]$ (200)
form a $(1-\alpha)$-CS for $\mu$ over $\mathcal{N}$.
Now, suppose $\{X_{n}\}\overset{\mathrm{iid}}{\sim}P$ with mean $\mu$ and variance $\sigma^{2}$. By Waudby-Smith et al. [2024, Lemma A.1], there exist $\{G_{n}\}\overset{\mathrm{iid}}{\sim}N_{0,1}$ such that
$\widehat{\mu}_{n}(X)-\mu={s_{n}(X)}\widehat{\mu}_{n}(G)+\varepsilon_{n},\quad\varepsilon_{n}=o(\sqrt{\log\log n/n}).$ (201)
Applying Theorem 4.9 to $\{G_{n}\}$, we have, with probability at least $1-\alpha$,
$|\widehat{\mu}_{n}(G)|\leqslant\rho(n,c,\alpha)s_{n}(G).$ (202)
Therefore, with probability at least $1-\alpha$,
$|\widehat{\mu}_{n}(X)-\mu|$ $\leqslant|\widehat{\mu}_{n}(X)-\mu-{s_{n}(X)}\widehat{\mu}_{n}(G)|+{s_{n}(X)}|\widehat{\mu}_{n}(G)|$ (203)
$\leqslant|\varepsilon_{n}|+{s_{n}(X)\rho(n,c,\alpha)s_{n}(G)}.$ (204)
Finally, we recall from our previous analysis that $s_{n}(X)\rho(n,c,\alpha)=\mathcal{O}(\sqrt{\log n/n})=\Omega(\varepsilon_{n})$, therefore
$\frac{|\varepsilon_{n}|+{s_{n}(X)\rho(n,c,\alpha)s_{n}(G)}}{s_{n}(X)\rho(n,c,\alpha)}=\frac{|\varepsilon_{n}|}{s_{n}(X)\rho(n,c,\alpha)}+{s_{n}(G)}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0+1=1,$ (205)
concluding that (200) is an $(1-\alpha)$-asymptotic CS over $\mathcal{L}^{2}$. ∎